# A minimalistic and general weighted averaging method for inconsistent data

### v1

@@ Abstract
The weighted average of inconsistent data is a common and tedious problem that many scientists have encountered. The standard weighted average is not recommended for these cases, and different alternative methods are proposed in the literature. Here, we introduce a new method based on Bayesian statistics for a broad application that keeps the number of assumptions to a minimum. The uncertainty associated with each input value is considered just a lower bound of the true unknown uncertainty. By assuming a non-informative (Jeffreys’) prior for true uncertainty and marginalising over its value, a modified Gaussian distribution is obtained with smoothly decreasing wings, which allows for a better treatment of scattered data and outliers. The proposed method is tested on a series of data sets: simulations, CODATA recommended value of the Newtonian gravitational constant, and some particle properties from the Particle Data Group, including the proton charge radius and the mass of the W boson. For the latter in particular, contrary to other works, our prediction lies in good agreement with the Standard Model. A freely available Python library is also provided for a simple implementation of our averaging method.
@@ I Introduction
where $n$ is the number of $x_{i}$ data points. In this way, the final value of the reduced $\chi^{2}$ is adjusted to be close to unity, as expected for consistent data sets. The use of the Birge ratio is indeed one of the pillars of the statistical treatment employed by the Task Group on Fundamental Constants of the Committee on Data of the International Science Council (here simply abbreviated by CODATA) and the Particle Data Group (PDG). A modified version of the Birge ratio has been proposed in past works based on Bayesian statistics. The scaling factor $R$ between the estimated uncertainty $\sigma_{i}$ and the real uncertainty $\sigma_{i}^{\prime}$ is considered unknown but common to all data points [3, 4, 5, 1]. Assuming a non-informative Jeffreys’ prior probability $p(R)\propto 1/R$ and marginalising over the possible values of $R$, a final probability distribution is obtained with an average value (the mode of the final probability distribution) equal to the standard weighted average, and an uncertainty that corresponding to an expansion of the standard weighted average uncertainty by a factor equal to $R_{\mathrm{Bayes}}=\sqrt{(n-1)/(n-3)}R_{\mathrm{Birge}}$. Variations of this approach are discussed in Refs. [6, 7].
@@
In principle, because of the common scaling factor for each measurement result, the Birge ratio and its modified versions discussed above are not well adapted to inter-laboratory averages, for which very different systematic effects can occur. To compensate partially for such an issue, past works [3, 7] proposed to assign a random bias $\beta_{i}$ to each measurement, with a common mean value and standard deviation $\sigma_{\mathrm{bias}}$ for the entire ensemble of measurements. Here, a double marginalisation over the $\beta_{i}$ values and their shared uncertainty $\sigma_{\mathrm{bias}}$ is required (for which a prior probability has to be chosen, generally a non-informative Jeffreys’ prior). An evolution of such an approach has been proposed [8]. It consists of organising the input values into clusters, each with a different $\sigma_{\mathrm{bias}}$ value, and subsequently performing a Bayesian model average, in which each model corresponds to a different clustering choice.
@@
An alternative approach that, like the Birge ratio, avoids formulating any hypothesis on the nature of missing systematic uncertainty consists of estimating the uncertainty directly from the data dispersion with no particular assumption about the associated probability distribution [3, 9]. Datum-by-datum, the associated uncertainty $\tilde{\sigma}_{i}$ is obtained by a quadratic sum $\tilde{\sigma}_{i}=\sqrt{\sigma_{i}^{2}+d_{i}^{2}}$ of the known uncertainty $\sigma_{i}$ and the estimated missing uncertainty obtained from the difference $d_{i}=x_{i}-\hat{\mu}$ between the input value $x_{i}$ and the standard weighted average itself. Since $\hat{\mu}$ is thus present in both the left and right expressions of Eq. (1), the final average is obtained recursively. Like the original formulation of the Birge ratio, this method lacks statistical foundations but has the benefit of being very simply formulated.
@@
All the previously described methods share the implicit assumption that the uncertainty $\sigma_{i}$ is a lower bound of the real uncertainty $\sigma_{i}^{\prime}$. This simple and clear statement has been translated into formulas by Sivia and Skilling in 2004 [10] avoiding common scaling factors (like $R_{\mathrm{Birge}}$) or random bias dispersion (like $\beta_{\mathrm{bias}}$), but considering, for each point, a modification of the Gaussian distribution by the marginalisation over $\sigma_{i}^{\prime}$. For this approach, a prior probability $p(\sigma_{i}^{\prime})$ for $\sigma_{i}^{\prime}$ has to be chosen. The natural choice would be the non-informative Jeffreys’ prior $p(\sigma_{i}^{\prime})\propto 1/\sigma_{i}^{\prime}$. If not constraint by an upper bound, this choice causes divergence because of the non-integrability of the resulting final probability. To avoid this problem, a modified prior $\propto 1/(\sigma_{i}^{\prime})^{2}$ has been proposed and discussed in Refs. [10, 11, 6]. Other more complex approaches with no lower bounds for $\sigma_{i}^{\prime}$ can be found in Refs. [10, 1].
Here, we consider the conservative approach with very few assumptions proposed by Sivia and Skilling [10]. Unlike Sivia and Skilling and other more recent works [10, 11, 6], we adopt a purely Jeffreys’ distribution for $\sigma^{\prime}_{i}$, taking some precautions to avoid possible divergences. This is obtained by studying the limit case of $\sigma^{\mathrm{max}}_{i}\to\infty$ with $\sigma^{\mathrm{max}}_{i}$ indicating the upper bound of $\sigma_{i}^{\prime}$. The final probability distribution associated with each datum $x_{i}$ is no longer Gaussian, implying significant modifications of the final weighted average. The consequences of such modifications are discussed and compared to other methods in Section III, using practical cases with simulated and real data. Details on the derivation of the method are presented in Sec. II. In section IV, the Python library based on the introduced method is presented. The final section is devoted to the conclusion.
@@ II.1 General considerations
The standard weighted average of independent measurements $x_{i}$ with uncertainties $\sigma_{i}$ is obtained by maximising the total probability of the mean value $\mu$ obtained by the product of the single probabilities for each (assumed independent) measurement, with
@@
$\sigma_{\hat{\mu}}=\left.\frac{\partial^{2}}{\partial\mu^{2}}\log[p(\mu|\{x_{i},\sigma_{i}\})]\right|_{\mu=\hat{\mu}}.$ (5)
In line with standard methods, we consider the average as the best value $\hat{\mu}$ of $\mu$ that maximises Eq. (4) (assuming the Jeffreys’ prior $p(\mu)=\text{const.}$.), with its uncertainty given by Eq. (5), but where the single probabilities $p(x_{i}|\mu,\sigma_{i})$ are no longer Gaussian, but are instead obtained by assuming certain hypotheses on the priors and performing marginalisations.
Our final goal is to provide the scientific community with a new tool to obtain a robust weighted average that can be easily understood and, more importantly, easily implemented as an alternative to the standard weighted average. Two prerogatives are thus essential: to propose something very general and to consider a minimal number of assumptions. The generality is particularly important to treat very common but different scenarios of inconsistent data averaging: i) from measurements obtained with the same apparatus (with a common uncontrolled systematic effect) or ii) from different types of measurements in different laboratories (with possible uncorrelated biases). For these purposes, we adopt a pessimistic framework where the uncertainties $\sigma_{i}$ are regarded as a lower bound of the real uncertainty $\sigma^{\prime}_{i}$, without any assumptions on the possible biases and relations influencing the available data set (similarly to Refs. [10, 14]). Any systematic error is considered to be included in the global and unknown uncertainty $\sigma_{i}^{\prime}$.
@@ II.2 Sivia and Skilling’s conservative weighted average
Compared with a Gaussian distribution, the above expression is characterised by a significantly larger spread, with tails proportional to $1/(x_{i})^{2}$. Once plugged into Eq. (4), such slowly descending tails supply sufficient flexibility to be tolerant of inconsistent data. Unlike the standard weighted average, the maximising value $\hat{\mu}$ and its associated uncertainty $\sigma_{\hat{\mu}}$ have no analytical form like the standard weighted average (Eqs. (1),(2)), but can be easily determined with numerical methods. Now, both $\hat{\mu}$ and $\sigma_{\hat{\mu}}$ depend on the data spread. Note that because of the presence of the tails, even for consistent data sets, the final uncertainty of the average is generally greater than the one obtained by the standard method.
@@ II.3 Limit solution with Jeffreys’ prior
A criticism that could be directed at Eq. (9) is that the choice of the prior probability distribution of $\sigma_{i}^{\prime}$ in Eq. (8) does not respect the non-informative criterion of Jeffreys’ prior. The solution proposed here to keep Jeffreys’ prior without introducing any new parameters is to consider the limit case $\sigma^{\mathrm{max}}_{i}\to\infty$. The divergence of the resulting probability distribution is circumvented by considering only its maximum and the second-order log-derivative for the estimation of the limit value of the weighted average and the associated uncertainty, respectively.
@@
Compared to Eq. (9), we can note that the distribution tails decrease even more smoothly than for the conservative approach, with a dependency of $1/x_{i}$ instead of $1/(x_{i})^{2}$ that can even better tolerate the presence of outliers. The logarithmic form of the total probability $p(\mu|\{x_{i},\sigma_{i},\sigma^{\mathrm{max}}_{i}\})$ is thus given by
@@
This constant term does not play a role in the search for $\hat{\mu}$, as it depends only on the $\sigma_{i}^{\prime}$ boundaries.
@@
where the constant $C^{\infty}=\lim_{\sigma^{\mathrm{max}}_{i}\to\infty}C=\infty$ is indeed divergent, but the position of the maximum and the value of the second derivative, and thus the weighted average and its uncertainty are still well defined. As in the case of the conservative weighted average, no analytical solution is available for $\hat{\mu}$ and $\sigma_{\hat{\mu}}$, so the solution must be found numerically.
@@ III Some applications
In this section, we present a series of applications for common data analysis cases. In the first subsection, we will study simulated data with known theoretical values of mean and standard deviation. In the second subsection, an analysis of the different values of the Newtonian gravitational constant from past CODATA compilations is proposed. The third subsection is dedicated to fundamental particle properties, including the controversial data sets of the W boson mass and the proton charge radius. Additional applications of the proposed method for high-resolution x-ray spectroscopy can be found in Ref. [15].
@@ 3. Set 3:
For the three sets, the standard weighted average, with or without Birge ratio correction, is compared to the conservative and Jeffreys’ weighted averages. The results are presented tables 13 and in Fig. 1 (for the standard and Jeffreys’ averages only). As we can see, for normal and inconsistent data without outliers, the mean value is well reproduced by all methods. Because of the pessimistic priors on $\sigma_{i}^{\prime}$, Jeffreys’ and conservative final uncertainties are generally larger than the standard one. For consistent data (set 1), they are larger by a factor of about two. As expected, for the inconsistent data (set 2), the uncertainty associated with the standard weighted average is significantly smaller than the others, with the Jeffreys’ uncertainty being the largest, followed by the conservative and the Birge uncertainties. Jeffreys’ average uncertainty is almost six time larger than the standard weighted average uncertainty and double that of the Birge-ratio-corrected uncertainty.
When an outlier is present, Jeffreys’ average values are quite different from the standard one. The effect of the presence of an outlier (set 3) is clearly visible in Fig. 2, where the final likelihoods are plotted together with the data, and in the results presented in in table 3. As we can see, the effect on the standard likelihood, obtained by the product of Gaussian distributions, is drastic, with a shift in the direction of the outlier. If the data uncertainties are regarded as lower bounds only, the effect is greatly mitigated, resulting in just an asymmetry of the tails for Jeffreys’ and conservative priors, which have very similar final probability distributions (once normalised). This behaviour looks quite similar to other methods that utilize the deviation of the data point from the calculated value $\hat{\mu}$ to estimate the possible missing uncertainty contributions [3, 9]. However, unlike these past works, here the distributions are derived from the initial assumptions on the uncertainties, for which we consider a priori that the provided values $\sigma_{i}$ are only lower bounds of the real uncertainty.
@@ III.2 The Newtonian constant of gravity
Figure 3: Comparison between the official CODATA values [18, 19, 20, 21, 22, 23, 24, 25, 26] of the Newtonian constant and the values obtained by the Bayesian weighted average using Jeffreys’ prior. CODATA values obtained from single measurements are presented alone, as no weighted average could be performed. The small error bar of the CODATA values indicates the uncertainty calculated by the standard weighted average, and the large one indicates the recommended uncertainty. The horizontal dashed line corresponds to the latest CODATA value (2022 edition [27], equal to the 2018 edition one).
Here, we apply our averaging method to all data sets included in the different editions of the CODATA compilation [18, 19, 20, 21, 22, 23, 24, 25, 26]. The results are presented together with the official values in Fig. 3. For each reported CODATA value, the large error bar corresponds to the recommended value of the uncertainty, and the small one to the uncertainty calculated by the standard weighted average. As we can see, the standard weighted average is, for some years, several standard deviations away from the most recent CODATA value from 2018 (the horizontal dashed line in the figure, which is the same as the more recent CODATA 2022 compilation [27]), considered here as the reference value. Contrary to standard procedures, one can see that the values obtained by the Jeffreys’ average are consistently in good agreement, being less than one standard deviation away from the most recent CODATA value, and are characterised by a more plausible uncertainty.
@@
The 1998 case is particularly difficult due to the inconsistency within the data set, arising from one very precise measurement [28] that differs significantly from the average of the other measurements. The value was later found to be affected by a large systematic error [29]. Details of the analysis of this specific case are presented in table 4 and Fig. 4. The CODATA recommended value is obtained by a standard weighted average of all values, excluding the suspicious measurement. The corresponding uncertainty is obtained by applying an expansion factor of 37 to the standard weighted average uncertainty to reflect the presence of the outlier. More precisely, the final uncertainty has been chosen to ensure that the difference between the recommended value and the outlier is four times larger than the final uncertainty. Like in set 3 of the previous section, the Jeffreys’ average is only slightly affected by the outlier in this challenging case.

### v2

@@ Abstract
The weighted average of inconsistent data is a common and tedious problem that many scientists have encountered. The standard weighted average is not recommended for these cases, and different alternative methods are proposed in the literature. Here, we discuss a method first proposed by Sivia in 1996 that is based on Bayesian statistics and keeps the number of assumptions to a minimum. We propose this approach as a new standard for calculating weighted averages. The uncertainty associated with each input value is considered to be just a lower bound of the true unknown uncertainty. The resulting likelihood function is no longer Gaussian, but has smoothly decreasing wings, which allows for a better treatment of scattered data and outliers. The proposed method is tested on a series of data sets: simulations, CODATA recommended value of the Newtonian gravitational constant, and some particle properties from the Particle Data Group, including the proton charge radius. A freely available Python library is also provided for a simple implementation of the proposed averaging method.
@@ I Introduction
where $n$ is the number of $x_{i}$ data points. In this way, the final value of the reduced $\chi^{2}$ is adjusted to be close to unity, as expected for consistent data sets. The use of the Birge ratio is indeed one of the pillars of the statistical treatment employed by the Task Group on Fundamental Constants of the Committee on Data of the International Science Council (here simply abbreviated by CODATA) and the Particle Data Group (PDG). A modified version of the Birge ratio has been proposed in past works based on Bayesian statistics. The scaling factor $R$ between the estimated uncertainty $\sigma_{i}$ and the real uncertainty $\sigma_{i}^{\prime}$ is considered unknown but common to all data points [3, 4, 5, 1]. Assuming a non-informative Jeffreys’ prior probability $p(R)\propto 1/R$ and marginalising over the possible values of $R$, a final probability distribution is obtained with an average value (the mode of the final probability distribution) equal to the standard weighted average. The uncertainty corresponds to an expansion of the standard weighted average uncertainty by a factor equal to $R_{\mathrm{Bayes}}=\sqrt{(n-1)/(n-3)}R_{\mathrm{Birge}}$. Variations of this approach are discussed in Refs. [6, 7].
@@
In principle, because of the common scaling factor for each measurement result, the Birge ratio and its modified versions discussed above are not well suited for inter-laboratory averages, where very different systematic effects can occur. To compensate partially for such an issue, past works [3, 7] proposed to assign a random bias $\beta_{i}$ to each measurement, with a common mean value and standard deviation $\sigma_{\mathrm{bias}}$ for the entire ensemble of measurements. Here, a double marginalisation over the $\beta_{i}$ values and their shared uncertainty $\sigma_{\mathrm{bias}}$ is required, for which a prior probability has to be chosen, typically a non-informative Jeffreys’ prior. An evolution of such an approach has been proposed [8]. It consists of organising the input values into clusters, each with a different $\sigma_{\mathrm{bias}}$ value, and subsequently performing a Bayesian model average, in which each model corresponds to a different clustering choice.
@@
An alternative approach, which like the Birge ratio avoids formulating any hypothesis on the nature of the missing systematic uncertainty, consists of estimating the uncertainty directly from the data dispersion with no particular assumption about the associated probability distribution [3, 9]. Datum-by-datum, the associated uncertainty $\tilde{\sigma}_{i}$ is obtained by a quadratic sum $\tilde{\sigma}_{i}=\sqrt{\sigma_{i}^{2}+d_{i}^{2}}$ of the known uncertainty $\sigma_{i}$ and the estimated missing uncertainty obtained from the difference $d_{i}=x_{i}-\hat{\mu}$ between the input value $x_{i}$ and the standard weighted average itself. Since $\hat{\mu}$ is thus present in both the left and right expressions of Eq. (1), the final average is obtained recursively. Like the original formulation of the Birge ratio, this method lacks statistical foundations but has the benefit of being very simply formulated.
@@
All the previously described methods share the implicit assumption that the uncertainty $\sigma_{i}$ is a lower bound of the real uncertainty $\sigma_{i}^{\prime}$. This simple and clear statement has been translated into formulas by Sivia and Skilling in 2004 [10], avoiding common scaling factors (like $R_{\mathrm{Birge}}$) or random bias dispersion (like $\beta_{\mathrm{bias}}$), but considering for each point a modification of the Gaussian distribution by the marginalisation over $\sigma_{i}^{\prime}$. For this approach, a prior probability $p(\sigma_{i}^{\prime})$ for $\sigma_{i}^{\prime}$ has to be chosen. The natural choice would be the non-informative Jeffreys’ prior $p(\sigma_{i}^{\prime})\propto 1/\sigma_{i}^{\prime}$. If not constraint by an upper bound, this choice causes divergence because of the non-integrability of the resulting final probability distribution associated to each datum. To avoid this problem, a modified prior $p(\sigma_{i}^{\prime})\propto 1/(\sigma_{i}^{\prime})^{2}$ has been proposed and discussed in Refs. [10, 11, 6]. Alternatively, the inverse-gamma distribution, which is integrable but introduces additional parameters, is proposed as a prior distribution [1, 12]. Other more complex approaches with no lower bounds for $\sigma_{i}^{\prime}$ can be found in Refs. [10, 1].
Another approach involves the assumption of a likelihood that is not Gaussian. For example, Dose [13] proposes using Laplace and hyperbolic cosine likelihoods in addition to the Gaussian likelihood and averaging the three distributions by applying Bayesian model selection.
Here, we consider an approach with very few assumptions, similar to the work of Sivia and Skilling [10], but adopt the Jeffreys’ non-informative prior, like in Sivia’s earlier work [14]. Our intent is not to propose an automatic approach to replace the work of data selection, but to offer an easily implementable method as a common reference for averaging, while avoiding the issues associated with the standard weighted average. For these reasons, like the standard weighted average, we keep the basic assumptions simple and very general, with the risk of being too conservative in the estimation of the final uncertainty in certain cases. To facilitate the application of the proposed method in daily data analysis, a Python code is provided.
Details on the derivation of the method are presented in Sec. II. In Section III, the proposed method is compared to the other methods in a series of cases using simulated and real data. In section IV, the Python library based on the introduced method is presented. The final section is devoted to the discussion and conclusion.
@@ II.1 General considerations
The standard weighted average of independent measurements $x_{i}$ with uncertainties $\sigma_{i}$ is obtained by maximising the total probability of the mean value $\mu$, which results from the product of the single probabilities for each (assumed independent) measurement, with
@@
$\sigma_{\hat{\mu}}=\left(-\frac{\partial^{2}}{\partial\mu^{2}}\log[p(\mu|\{x_{i},\sigma_{i}\})]\bigg{|}_{\mu=\hat{\mu}}\right)^{-\frac{1}{2}}.$ (5)
In line with standard methods, we consider the average as the value $\hat{\mu}$ that maximises Eq. (4) (assuming the Jeffreys’ prior $p(\mu)=\text{const.}$.), with its uncertainty given by Eq. (5). The single probabilities $p(x_{i}|\mu,\sigma_{i})$ are no longer Gaussian, but are instead obtained by assuming certain hypotheses on the priors and performing marginalisations.
As wrote above, our final goal is to provide the scientific community with a new accessible tool to obtain a robust weighted average that can be easily understood. Two prerogatives are thus essential: to propose something very general and to consider a minimal number of assumptions. The generality is particularly important to treat very common but different scenarios of inconsistent data averaging: i) from measurements obtained with the same apparatus (with a common uncontrolled systematic effect) or ii) from different types of measurements in different laboratories (with possible uncorrelated biases). For these purposes, we adopt a pessimistic framework where the uncertainty $\sigma_{i}$ is regarded as a lower bound of the real uncertainty $\sigma^{\prime}_{i}$, without any additional assumptions on the possible biases and relations influencing the available data set (similarly to Refs. [10, 17]). Any systematic error is considered to be included in the unknown uncertainty $\sigma_{i}^{\prime}$.
@@ II.2 Sivia and Skilling’s conservative weighted average
Compared with a Gaussian distribution, as visible in Fig. 1, the above expression is characterised by a significantly larger spread, with tails proportional to $1/(x_{i})^{2}$. Once plugged into Eq. (4), such slowly descending tails supply sufficient flexibility to be tolerant of inconsistent data. Unlike the standard weighted average, the maximising value $\hat{\mu}$ and its associated uncertainty $\sigma_{\hat{\mu}}$ have no analytical form, but can be easily determined with numerical methods. Now, both $\hat{\mu}$ and $\sigma_{\hat{\mu}}$ depend on the data spread. Note that, because of the presence of the tails, even for consistent data sets, the final uncertainty of the weighted average is generally greater than the one obtained by the standard method.
Figure 1: Comparison between the different assumed probability distribution for each datum for $\mu=0,\sigma_{i}=1$.
@@ II.3 Limit solution with Jeffreys’ prior
A criticism that could be directed at Eq. (9) is that the choice of the prior probability distribution of $\sigma_{i}^{\prime}$ in Eq. (8) does not respect the non-informative criterion of Jeffreys’ prior. A possible solution to keep Jeffreys’ prior without introducing any new parameters is given by Sivia in Ref. [14]. It considers the limit case $\sigma^{\mathrm{max}}_{i}\to\infty$. The divergence of the resulting probability distribution is circumvented by considering only its maximum and the second-order log-derivative for the estimation of the limit value of the weighted average and the associated uncertainty, respectively.
@@
Compared to Eq. (9), we can note that the distribution tails decrease even more smoothly than in the conservative approach, with a dependency of $1/x_{i}$ instead of $1/x_{i}^{2}$, which can tolerate the presence of outliers even better. These more pronounced tails are well visible in Fig. 1. The logarithmic form of the total probability $p(\mu|\{x_{i},\sigma_{i},\sigma^{\mathrm{max}}_{i}\})$ is thus given by
@@
This constant term does not play a role in the search for $\hat{\mu}$, as $C$ depends only on the boundary values $\sigma_{i}$ and $\sigma_{i}^{\mathrm{max}}$.
@@
where the constant $C^{\infty}=\lim_{\sigma^{\mathrm{max}}_{i}\to\infty}C=\infty$ is indeed divergent, but for more than one data point the distribution is integrable, and the variance is finite for more than two data points. In particular, the position of the maximum and the value of the second derivative and thus the weighted average and its uncertainty are well defined. As in the case of the conservative weighted average, no analytical solution is available for $\hat{\mu}$ and $\sigma_{\hat{\mu}}$, so the solution must be found numerically.
We consider the formulas presented in this section as a new reference method for the weighted average. In fact, the conservative approach of Sec. II.2 still involves some arbitrariness in the choice of the prior for the average. This is not the case for the above formulas, where the universal choice of the Jeffreys’ prior is adopted. In the following section, we will refer to this approach as the Jeffreys’ weighted average.
@@ III Some applications
In this section, we present a series of applications for common data analysis cases. In the first subsection, we will study simulated data with known theoretical values of mean and standard deviation. In the second subsection, an analysis of the different values of the Newtonian gravitational constant from past CODATA compilations is proposed. The third subsection is dedicated to fundamental particle properties, including the controversial data set of the proton charge radius. Additional calculations of the Jeffreys’ weighted average for inconsistent data measured using the same experimental apparatus can be found in Ref. [18].
@@ III.1 Synthetic tests
For the three sets, the standard weighted average, with or without Birge ratio correction, is compared to the conservative and Jeffreys’ weighted averages. The results are presented in Tables 13 and in Fig. 2 (for the standard and Jeffreys’ weighted averages only) and Fig. 3. As we can see, for normal and inconsistent data without outliers, the mean value is well reproduced by all methods. Because of the pessimistic priors on $\sigma_{i}^{\prime}$, the Jeffreys’ and conservative final uncertainties are generally larger than the standard uncertainty. For consistent data (set 1), they are larger by a factor of about two. As expected, for the inconsistent data (set 2), the uncertainty associated with the standard weighted average is significantly smaller than the others, with the Jeffreys’ uncertainty being the largest, followed by the conservative and the Birge uncertainties. Jeffreys’ weighted average uncertainty is almost six time larger than the standard weighted average uncertainty and double that of the Birge-ratio-corrected uncertainty.
When an outlier is present, Jeffreys’ weighted average values are quite different from the standard weighted average. The effect of the presence of an outlier (set 3) is clearly visible in Fig. 3, where the final likelihoods are plotted together with the data, and in the results presented in in Table 3. As we can see, the effect on the standard likelihood, obtained by the product of Gaussian distributions, is drastic, with a shift in the direction of the outlier. If the data uncertainties are regarded as lower bounds only, the effect is greatly mitigated, resulting in just an asymmetry of the tails for the Jeffreys’ and conservative priors, which have very similar final probability distributions. This behaviour looks quite similar to other methods that utilize the deviation of the data point from the calculated value $\hat{\mu}$ to estimate the possible missing uncertainty contributions [3, 9]. However, unlike these past works, here the distributions are derived from the initial assumptions on the uncertainties, for which we consider a priori that the provided values $\sigma_{i}$ are only lower bounds of the real uncertainty.
@@ III.2 The Newtonian constant of gravity
Figure 4: Comparison between the official CODATA values [21, 22, 23, 24, 25, 26, 27, 28, 29] of the Newtonian constant and the values obtained by the Bayesian weighted average using Jeffreys’ prior. CODATA values obtained from single measurements are presented alone, as no weighted average could be performed. The small error bar of the CODATA values indicates the uncertainty calculated by the standard weighted average, and the large one indicates the recommended uncertainty. The horizontal dashed line corresponds to the latest CODATA value (2022 edition [30], equal to the 2018 edition value).
Here, we apply the proposed averaging method to all data sets included in the different editions of the CODATA compilation [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The results are presented together with the official values in Fig. 4. For each reported CODATA value, the large error bar corresponds to the recommended value of the uncertainty, and the small one to the uncertainty calculated by the standard weighted average. As we can see, the standard weighted average is, for some years, several standard deviations away from the most recent CODATA value from 2018 (the horizontal dashed line in the figure, which is the same as the more recent CODATA 2022 value [30]), considered here as the reference value. Contrary to standard procedures, one can see that the values obtained by the Jeffreys’ weighted average are consistently in good agreement, being less than one standard deviation away from the most recent CODATA value, and are characterised by a more plausible uncertainty.
@@
The 1998 case is particularly difficult due to the inconsistency within the data set, arising from one very precise measurement [31] that differs significantly from the average of the other measurements. The value was later found to be affected by a large systematic error [32]. Details of the analysis of this specific case are presented in Table 4 and Fig. 5. The CODATA recommended value is obtained by the standard weighted average of all values, excluding the suspicious measurement. The corresponding uncertainty is obtained by applying an expansion factor of 37 to the standard weighted average uncertainty to reflect the presence of the outlier. More precisely, the final uncertainty has been chosen to ensure that the difference between the recommended value and the outlier is four times larger than the final uncertainty. Like in set 3 of the previous section, the Jeffreys’ weighted average is only slightly affected by the outlier in this challenging case. Looking at Table 4, one can note that the Jeffreys’ weighted average has a significantly smaller uncertainty than the method tested on the same data set by Dose [13]. There, a Bayesian model average was applied on different likelihood functions.